Determine the mass of N2H4 obtained from 140 g of N2, according to the reaction: N2 + 2 H2 → N2H4.


Data: H-1, N-14.


1
Expert's answer
2021-06-21T05:29:05-0400

mass of "N_2H_4=?"

mass of "N_2=140g"

number of moles of nitrogen"=\frac{140}{2\times14}=5moles"

mole ratio is "1:1"

number of moles of "N_2H_4=5moles"

mass of "N_2H_4=5(2\times14+1\times4)=5\times32=160g"